# FedMAC: Tackling Partial-Modality Missing in Federated Learning with Cross-Modal Aggregation and Contrastive Regularization

### v1

@@ Abstract
Federated Learning (FL) is a method for training machine learning models using distributed data sources. It ensures privacy by allowing clients to collaboratively learn a shared global model while storing their data locally. However, a significant challenge arises when dealing with missing modalities in clients’ datasets, where certain features or modalities are unavailable or incomplete, leading to heterogeneous data distribution. While previous studies have addressed the issue of complete-modality missing [^1] Complete missing is when one or more modalities are absent in server and clients’ data., they fail to tackle partial-modality missing [^2] Partial missing is when only parts of one or more modalities are absent in server and clients’ data. on account of severe heterogeneity among clients at an instance level, where the pattern of missing data can vary significantly from one sample to another. To tackle this challenge, this study proposes a novel framework named FedMAC, designed to address multi-modality missing under conditions of partial-modality missing in FL. Additionally, to avoid trivial aggregation of multi-modal features, we introduce contrastive-based regularization to impose additional constraints on the latent representation space. The experimental results demonstrate the effectiveness of FedMAC across various client configurations with statistical heterogeneity, outperforming baseline methods by up to 26% in severe missing scenarios, highlighting its potential as a solution for the challenge of partially missing modalities in federated systems.
Index Terms:
Contrastive Learning, Federated Learning, Modality Missing, Multi-modal Learning.
@@ I Introduction
Multi-modal Federated Learning and Modality Missing Issue. Federated Learning (FL) [1] is a decentralized machine learning paradigm that allows multiple devices to train a shared model collaboratively while keeping data localized. This approach preserves privacy and reduces data transfer overhead, driven by growing concerns over data privacy and security [2] and the inefficiencies associated with transferring large volumes of data to central servers [3]. Traditionally, FL methods [4, 5, 6, 7, 8, 9, 10, 11] have primarily focused on single-modality data (e.g., image or text), which limits their scope of applications. With the rapid development of mobile phones and Internet of Things (IoT) devices [12], there is increasing potential for leveraging multi-modal data. Multi-modal learning integrates various types of data (e.g., sensory, visual, audio), providing a richer and more comprehensive understanding [13]. Combined with the privacy protection capabilities of FL, multi-modal FL enhances the performance and robustness of the system [14].
@@
In the past, multi-modal FL methods [14, 15, 16] assume that all modalities are sufficient and complete for every sample. This assumption limits the application of FL, as embedded sensors in mobile phones or IoT devices can fail or have poor contact, leading to missing modalities. The modality missing in FL can be classified into two types: complete missing and partial missing. The former refers to a scenario where different combinations of modalities owned by different clients, but the modalities must be complete, while the latter allows diverse modality incompleteness within clients’ datasets at an instance level. The missing modalities pose significant challenges to multi-modal learning in federated systems. As shown in Fig. 1, models trained on full modalities experience significant performance degradation when encountering missing modality data in the testing phase.
Existing Approaches in Dealing with Modality Missing Problems. In conventional centralized learning systems, there are two typical approaches in dealing with modality missing issues: imputation and non-imputation methods. The core concept of the imputation-based method  [17, 18] is to replace the missing modalities with synthetic data, which can be either zeros or the average values of the other modalities. However, these methods rely solely on intra-modal information and ignore inter-modal correlations, resulting in suboptimal performance. On the other hand, non-imputation methods [19, 20, 21] such as hypergraph-based and optimal sparse linear prediction-based approaches, aim to simultaneously benefit models from missing and existing modalities without adding artificial noises into training data. However, these methods assume centralized training, making them difficult to apply to FL, which relies on decentralized training to ensure privacy.
Despite the critical need for methods to address missing modalities in FL, research in this area remains limited. Chen et al. [22] and Phung et al. [23] have made early attempts to address this issue in scenarios involving the complete missing of one or multiple modalities, as shown in Fig. 1(b). However, as shown in Fig. 1(c), partial-modality missing is more common due to data collection issues or device malfunctions within the network. Partial-modality missing is more challenging than complete-modality missing, as the server cannot ascertain which devices have missing modalities, resulting in significant noise and fluctuations during training. Recently, Yu et al. [24] proposed FedInMM to address partial-modality missing. However, FedInMM assumes that some time stamps of each missing timeseries-based modality are preserved and employs a Long Short-Term Memory (LSTM)-based module to extract temporal information. This assumption limits its generalizability across different scenarios where temporal correlations may be weak or absent.
@@
  • We conduct extensive experiments across multiple scenarios of missing modalities to demonstrate the effectiveness of the proposed method in real-world conditions. The experimental results show the superiority of the proposed method compared to state-of-the-art methods on the subset of multi-modal PTB-XL dataset [25].
@@ II-B Overview of FedMAC
This study focuses on the local training phase and employs the widely-used FedAvg algorithm [1], which relies on weighted averaging for global model aggregation. A key challenge in multi-modal FL, where data types or “modalities” vary across clients, is modality heterogeneity. In simple terms, different clients may have access to distinct sets of modalities and observe incomplete collection of training samples, which can introduce noise and degrade model performance.
@@ II-C2 Cross-Modal Aggregator
$S_{Z}\in\mathbb{R}^{T_{k}\times T_{k}};\quad S^{u,v}_{Z}=\frac{u\cdot v^{T}}{||u||\cdot||v||};\quad u,v\in Z,$ (5)
@@ II-E Aggregation Operation on Server
In this study, FedMAC employs the FedAvg algorithm [1] for server aggregation to obtain the global model after each communication round. The server initializes the global model $\omega^{0}$, and after each round, clients replace their local models with the updated global model. Clients then update their models on local data for a few epochs before sending them back to the server. The server aggregates the model as follows:
@@ III Evaluation
Model Missing statistics $(p_{m}/p_{s})$ in the IID setting Missing statistics $(p_{m}/p_{s})$ in the Non-IID setting
@@ III-A Experimental Settings
Dataset and Preparation. Following Phung et al. [23], we use a subset of the PTB-XL dataset [25]. This subset consists of 3,963 clinical samples across five distinct classes. Each sample has 12 modalities (the electrocardiogram recordings) and is assigned to a single class. The dataset is subsequently divided into clients’ training datasets and a server dataset with the 80:20 ratio. The clients’ dataset is then divided among $K$ clients according to the IID or Non-IID settings.
@@
Baseline Methods. We compare the proposed FedMAC with several baseline methods, including FedProx [6], FedMSplit [22], MIFL [23], and FedInMM [24]. FedProx is a naive approach that does not explicitly address missing modalities, treating all instances equally regardless of their available modalities. FedMSplit and MIFL are methods designed to handle scenarios with complete-modality missing. On the other hand, FedInMM is the recently developed method for handling partial-modality missing scenarios.
@@
Implementation Details. We implement FedMAC as detailed in Sec. II, utilizing an Inception Network as the Modality-Specific Extractor, as following [23]. For the classification task, we use Cross Entropy Loss for $\mathcal{L}_{task}$. The embedding dimension is set to $d_{H}=128$. The number of clients is $K=32$, as following [24], with local training epochs $E=3$, a batch size $B=32$, and a learning rate of $\eta=0.01$ for the IID setting and $\eta=0.008$ for the Non-IID setting. The optimizer used is Stochastic Gradient Descent (SGD) [26]. Communication with the server occurs over $T=1000$ rounds. The parameter $\lambda$ in Eq. 11 is set to $0.1$ for $p_{m}\times p_{k}\leq 0.5$ and $0.2$ otherwise. The temperature hyperparameter in Eqs. 610, and 9 are set to $1$. For the other methods, we employ the configurations as specified in their original papers.
@@ III-B1 Similar missing statistics between the client and server
IID Setting Results. Fig. 6 and Table I present the results of the proposed FedMAC compared to other methods. The proposed method consistently outperforms the others across all experimented scenarios with various missing statistics. When the missing degree is low (e.g., missing statistic $p_{m}/p_{s}=1.0/0.1$, resulting in a missing degree of $p_{m}\times p_{s}=0.1$), the performance differences between FedMAC and the other methods are minimal, with all achieving similar accuracy levels. However, as the missing degree increases (e.g., $p_{m}\times p_{s}=0.5$ and $p_{m}\times p_{s}=0.8$), FedMAC demonstrates a significant advantage, maintaining higher accuracy than its counterparts. This trend becomes even more pronounced in more complex scenarios where the missing statistics of the clients are intricate, such as $/0.5$ 0.5, $0.8/0.5$, and $/0.8$ 0.8. While other methods experience more significant performance degradation, FedMAC remains robust and consistently achieves the highest accuracy in every case. These results underscore the ability of FedMAC to effectively manage modality heterogeneity, particularly as the missing degree increases in the IID setting.
@@
Model Server’s missing statistics $(p_{m}/p_{s})$
@@
Client’s missing statistics $(p_{m}/p_{s})$ 1.0/0.1 FedProx 76.54 68.72 51.83 70.49 51.83 76.54
@@
Non-IID Setting Results. As shown in Fig. 6 and Table I, FedMAC outperforms all other methods. While other methods experience significant drops in accuracy as the missing degree increases, the accuracy of FedMAC remains consistently higher. In challenging settings with high missing degrees, such as $1.0/0.8$, $0.8/0.5$, and $/0.8$ 0.8, FedMAC maintains a clear advantage, demonstrating its robustness and adaptability to non-IID data distributions. However, in the non-IID setting, each client’s data represents a different distribution, leading to inconsistencies in the accuracy of FedMAC when aggregating model updates at the server during each communication round. The reason for this is that we use the FedAvg algorithm [1] for server aggregation in these experiments, which struggles in the non-IID setting because it aggregates based on the number of client samples. In future work, we are considering more robust aggregation strategies to better handle the variability in client data distributions and improve overall model performance in non-IID scenarios.
@@ III-B2 Different missing statistics among client and server
Model Missing statistics $(p_{m}/p_{s})$

### v2

@@ Abstract
Federated Learning (FL) is a method for training machine learning models using distributed data sources. It ensures privacy by allowing clients to collaboratively learn a shared global model while storing their data locally. However, a significant challenge arises when dealing with missing modalities in clients’ datasets, where certain features or modalities are unavailable or incomplete, leading to heterogeneous data distribution. While previous studies have tried addressing the issue of complete-modality missing [^1] Inter-client missing is when one or more modalities are absent in server and clients’ data. and intra-client missing [^2] Partial missing is when only parts of one or more modalities are absent in server and clients’ data., they fail to remains generalization in To tackle this challenge, this study proposes a novel framework named FedMAC, designed to address multi-modality missing under conditions of partial-modality missing in FL. Additionally, to avoid trivial aggregation of multi-modal features, we introduce contrastive-based regularization to impose additional constraints on the latent representation space. The experimental results demonstrate the effectiveness of FedMAC across various client configurations with statistical heterogeneity, outperforming baseline methods by up to 26% in severe missing scenarios, highlighting its potential as a solution for the challenge of partially missing modalities in federated systems. Our source code is provided at https://github.com/nmduonggg/PEPSY
@@ I Introduction
Multi-modal Federated Learning and Modality Missing Issue. Federated Learning (FL) [mcmahan2017fedavg] is a decentralized machine learning paradigm that allows multiple devices to train a shared model collaboratively while keeping data localized. This approach preserves privacy and reduces data transfer overhead, driven by growing concerns over data privacy and security [zhang2022data] and the inefficiencies associated with transferring large volumes of data to central servers [wang2022communication]. Traditionally, FL methods [li2019convergence, t2020personalized, li2020federated, fallah2020personalized, cho2022heterogeneous, pham2023sem, nguyen2023cadis, nguyen2022feddrl] have primarily focused on single-modality data (e.g., image or text), which limits their scope of applications. With the rapid development of mobile phones and Internet of Things (IoT) devices [brunete2021smart], there is increasing potential for leveraging multi-modal data. Multi-modal learning integrates various types of data (e.g., sensory, visual, audio), providing a richer and more comprehensive understanding [kaur2021comparative]. Combined with the privacy protection capabilities of FL, multi-modal FL enhances the performance and robustness of the system [xiong2022unified].
@@
In the past, multi-modal FL methods [xiong2022unified, chen2024feddat, peng2024fedmm] assume that all modalities are sufficient and complete for every sample. This assumption limits the application of FL, as embedded sensors in mobile phones or IoT devices can fail or have poor contact, leading to missing modalities. The modality missing in FL can be classified into two types: complete missing and partial missing. The former refers to a scenario where different combinations of modalities owned by different clients, but the modalities must be complete, while the latter allows diverse modality incompleteness within clients’ datasets at an instance level. The missing modalities pose significant challenges to multi-modal learning in federated systems. As shown in Fig. 1, models trained on full modalities experience significant performance degradation when encountering missing modality data in the testing phase.
Existing Approaches in Dealing with Modality Missing Problems. In conventional centralized learning systems, there are two typical approaches in dealing with modality missing issues: imputation and non-imputation methods. The core concept of the imputation-based method  [zhang2020deep, zhou2022missing] is to replace the missing modalities with synthetic data, which can be either zeros or the average values of the other modalities. However, these methods rely solely on intra-modal information and ignore inter-modal correlations, resulting in suboptimal performance. On the other hand, non-imputation methods [yu2020optimal, chen2020hgmf, poklukar2022geometric] such as hypergraph-based and optimal sparse linear prediction-based approaches, aim to simultaneously benefit models from missing and existing modalities without adding artificial noises into training data. However, these methods assume centralized training, making them difficult to apply to FL, which relies on decentralized training to ensure privacy.
Despite the critical need for methods to address missing modalities in FL, research in this area remains limited. Chen et al. [chen2022fedmsplit] and Phung et al. [phung2024mifl] have made early attempts to address this issue in scenarios involving the complete missing of one or multiple modalities, as shown in Fig. 1(b). However, as shown in Fig. 1(c), partial-modality missing is more common due to data collection issues or device malfunctions within the network. Partial-modality missing is more challenging than complete-modality missing, as the server cannot ascertain which devices have missing modalities, resulting in significant noise and fluctuations during training. Recently, Yu et al. [yu2024fedinmm] proposed FedInMM to address partial-modality missing. However, FedInMM assumes that some time stamps of each missing timeseries-based modality are preserved and employs a Long Short-Term Memory (LSTM)-based module to extract temporal information. This assumption limits its generalizability across different scenarios where temporal correlations may be weak or absent.
@@
  • We conduct extensive experiments across multiple scenarios of missing modalities to demonstrate the effectiveness of the proposed method in real-world conditions. The experimental results show the superiority of the proposed method compared to state-of-the-art methods on the subset of multi-modal PTB-XL dataset [dataset].
@@ II-B Overview of FedMAC
This study focuses on the local training phase and employs the widely-used FedAvg algorithm [mcmahan2017fedavg], which relies on weighted averaging for global model aggregation. A key challenge in multi-modal FL, where data types or “modalities” vary across clients, is modality heterogeneity. In simple terms, different clients may have access to distinct sets of modalities and observe incomplete collection of training samples, which can introduce noise and degrade model performance.
@@ II-C2 Cross-Modal Aggregator
[…]
@@ II-E Aggregation Operation on Server
In this study, FedMAC employs the FedAvg algorithm [mcmahan2017fedavg] for server aggregation to obtain the global model after each communication round. The server initializes the global model $\omega^{0}$, and after each round, clients replace their local models with the updated global model. Clients then update their models on local data for a few epochs before sending them back to the server. The server aggregates the model as follows:
@@ III Evaluation
Model
Missing statistics $(p_{m}/p_{s})$ in the IID setting Missing statistics $(p_{m}/p_{s})$ in the Non-IID setting
@@ III-A Experimental Settings
Dataset and Preparation. Following Phung et al. [phung2024mifl], we use a subset of the PTB-XL dataset [dataset]. This subset consists of 3,963 clinical samples across five distinct classes. Each sample has 12 modalities (the electrocardiogram recordings) and is assigned to a single class. The dataset is subsequently divided into clients’ training datasets and a server dataset with the 80:20 ratio. The clients’ dataset is then divided among $K$ clients according to the IID or Non-IID settings.
@@
Baseline Methods. We compare the proposed FedMAC with several baseline methods, including FedProx [li2020federated], FedMSplit [chen2022fedmsplit], MIFL [phung2024mifl], and FedInMM [yu2024fedinmm]. FedProx is a naive approach that does not explicitly address missing modalities, treating all instances equally regardless of their available modalities. FedMSplit and MIFL are methods designed to handle scenarios with complete-modality missing. On the other hand, FedInMM is the recently developed method for handling partial-modality missing scenarios.
@@
Implementation Details. We implement FedMAC as detailed in Sec. II, utilizing an Inception Network as the Modality-Specific Extractor, as following [phung2024mifl]. For the classification task, we use Cross Entropy Loss for $\mathcal{L}_{task}$. The embedding dimension is set to $d_{H}=128$. The number of clients is $K=32$, as following [yu2024fedinmm], with local training epochs $E=3$, a batch size $B=32$, and a learning rate of $\eta=0.01$ for the IID setting and $\eta=0.008$ for the Non-IID setting. The optimizer used is Stochastic Gradient Descent (SGD) [ruder2016overview]. Communication with the server occurs over $T=1000$ rounds. The parameter $\lambda$ in Eq. 11 is set to $0.1$ for $p_{m}\times p_{k}\leq 0.5$ and $0.2$ otherwise. The temperature hyperparameter in Eqs. 610, and 9 are set to $1$. For the other methods, we employ the configurations as specified in their original papers.
@@ III-B1 Similar missing statistics between the client and server
IID Setting Results. Fig. 6 and Table I present the results of the proposed FedMAC compared to other methods. The proposed method consistently outperforms the others across all experimented scenarios with various missing statistics. When the missing degree is low (e.g., missing statistic $p_{m}/p_{s}=1.0/0.1$, resulting in a missing degree of $p_{m}\times p_{s}=0.1$), the performance differences between FedMAC and the other methods are minimal, with all achieving similar accuracy levels. However, as the missing degree increases (e.g., $p_{m}\times p_{s}=0.5$ and $p_{m}\times p_{s}=0.8$), FedMAC demonstrates a significant advantage, maintaining higher accuracy than its counterparts. This trend becomes even more pronounced in more complex scenarios where the missing statistics of the clients are intricate, such as $0.5/0.5$, $0.8/0.5$, and $0.8/0.8$. While other methods experience more significant performance degradation, FedMAC remains robust and consistently achieves the highest accuracy in every case. These results underscore the ability of FedMAC to effectively manage modality heterogeneity, particularly as the missing degree increases in the IID setting.
@@
Model
Server’s missing statistics $(p_{m}/p_{s})$
@@
Client’s missing statistics $(p_{m}/p_{s})$
1.0/0.1
FedProx 76.54 68.72 51.83 70.49 51.83 76.54
@@
Non-IID Setting Results. As shown in Fig. 6 and Table I, FedMAC outperforms all other methods. While other methods experience significant drops in accuracy as the missing degree increases, the accuracy of FedMAC remains consistently higher. In challenging settings with high missing degrees, such as $1.0/0.8$, $0.8/0.5$, and $0.8/0.8$, FedMAC maintains a clear advantage, demonstrating its robustness and adaptability to non-IID data distributions. However, in the non-IID setting, each client’s data represents a different distribution, leading to inconsistencies in the accuracy of FedMAC when aggregating model updates at the server during each communication round. The reason for this is that we use the FedAvg algorithm [mcmahan2017fedavg] for server aggregation in these experiments, which struggles in the non-IID setting because it aggregates based on the number of client samples. In future work, we are considering more robust aggregation strategies to better handle the variability in client data distributions and improve overall model performance in non-IID scenarios.
@@ III-B2 Different missing statistics among client and server
Model
Missing statistics $(p_{m}/p_{s})$